# Towards Clean-Label Backdoor Attacks in the Physical World

### v1

@@ 5 Accidental backdoor activations
For CLPBA to be stealthy, we argue that it needs to be source-specific and trigger-specific. That is, (1) the trigger should be discriminative so that the victim model does not incorrectly recognize similar-looking objects as triggers, and (2) only the source class can use the trigger to manipulate the victim model’s prediction at inference time. We introduce two new metrics, False Positive Rate and Suspicion Rate, to test the specificity (and also stealthiness) of CLPBA:
@@
Table 5: Measurement of FPR and SR for GM attack ($\alpha=0.1,\epsilon=16/226$). A detailed set-up of our dataset to evaluate SR and FPR is given in Appendix C. We note that the big difference in FPR between fake beard and sunglasses triggers is due to selection bias in our choice of false triggers.
Poison pair Trigger Loss type ASR (%) ACC (%) FPR $\downarrow$ (%) SR $\downarrow$ (%)
9-5 Fake beard Original 98.7 (0.0) 99.9 (0.1) 50.8 (1.9) 12.5 (1.0)
With repel 98.7 (0.0) 99.9 (0.1) 49.7 (2.2) 3.9 (0.1)
7-3 Sunglasses Original 97.5 (1.1) 99.3 (0.1) 0.7 (0.8) 13.9 (0.7)
With repel 97.5 (1.4) 99.4 (0.2) 0.9 (0.1) 2.5 (0.3)
@@
$L_{adv}=N\cdot\overbrace{\sum_{x\in\widetilde{D}_{s}}\mathcal{L}\bigl{(}F_{\theta}(x),y_{t}\bigr{)}}^{\text{adversarial loss}}+\sum_{i\neq s}^{N}\overbrace{\Bigl{(}\sum_{x\in\widetilde{D}_{i}}\mathcal{L}\bigl{(}F_{\theta}(x),y_{i}\bigr{)}-\mathcal{L}\bigl{(}F_{\theta}(x),y_{t}\bigr{)}\Bigr{)}}^{\text{repel term}}$ (12)
@@ 6 Conclusion
Through extensive experiments, our research underscores the vulnerability of DNNs to clean-label backdoor attacks in the physical world. We thoroughly analyze the influence of poisoning algorithms, physical triggers, and source-target class pairs on the effectiveness of these attacks. Additionally, we propose novel pixel and feature regularizations to improve the imperceptibility of attacks in both pixel and feature spaces. A significant limitation identified is accidental backdoor activations of the poisoned model due to false triggers or non-source classes, which could compromise attack stealthiness. To address this, we propose incorporating repel terms into the adversarial objective to enhance trigger specificity to the source class. Our findings highlight the critical need for robust defenses against clean-label physical backdoor attacks and pave the way for future research on improving the resilience of DNNs against such threats.
Appendix A Procedure of blended label-consistent attack
Appendix B Natural physical backdoors in a dataset
@@
(a) Clean samples (without trigger)
@@
Appendix C Dataset
C.1 Ethics and data protection
IRB approval. Before conducting our study, we submitted a "Human Subjects Research – Expedited Review Form" to our country’s Institutional Review Board (IRB). Our study received approval from the chairman of the institutional ethical review board under decisions No 24/2016/QD-VINMEC, No 23/2016/QD-VINMEC, and No 77/2021/QD-VINMEC. We prepared a consent form beforehand to ensure transparency in the procedure of dataset collection. All 10 volunteers in our dataset provided explicit written consent for us to take their photos.
@@ C.4 Dataset set-up
False triggers. To evaluate the sensitivity of poisoned models to false triggers, as measured by false positive rate (FPR) (see Appendix D.4), we select three additional false triggers for each of the main triggers. Because it is hard to find a representative set of false triggers for each of the main triggers, we select the false triggers by changing the sizes, colors, and shapes of the objects (e.g, we change the shape of earrings and hats, change the colors of the headband, change the size of sticker). We also select objects that have similar locations on the face as false triggers (e.g., headbands are selected as false triggers for the hat trigger and the sunglasses trigger). In Figure 11, each row corresponds to a main trigger on the left, followed by three false triggers on the right.
@@
Appendix D Experiment settings
D.1 Computational resources.
All the experiments are conducted on the two servers. The first one with 7 RTX 3090 Ti 24 GB GPUS, and the second one contains 6 RTX A5000 24 GB GPUS. The code is implemented in Pytorch. We develop the framework based on the previous works [21, 9].
D.2 Training settings.
We evaluate the attacks on four Convolutional Neural Networks for facial recognition: ResNet50 [26], VGG16 [41], MobileNetV4 [42], Inception-ResNet [43]. We choose fine-tuning as the training method for the victim model since it is a widely adopted training method for its speed of convergence and accuracy. Compared to transfer-learning, Fine-tuning and training from scratch are shown to be hard scenarios for clean-label attacks [20, 21, 9]. Our models are pre-trained on VGGFace2 [27], a large-scale facial dataset containing 3.31 million images divided into 9131 classes. Since BLC requires an adversarially trained model, we perform fast adversarial training [44] for the pre-trained model on our dataset before crafting poisons. We do not apply any data augmentations to the training pipeline. We only standardize the training data using z-score normalization with the mean and standard variation calculated from the data. We observe that the training models converge very quickly and achieve the validation accuracy of over $99\%$ by epoch 10.
Table 8: Hyperapaters
@@ Appendix D Experiment settings
D.3 Attack settings.
Our experiments follow the gray-box scenario in which the attacker model has the same architecture and is trained on the same dataset as the victim model but has different weight initialization, and the attacker cannot access and modify the weights and training procedure of the victim model. Since different poisoning algorithms have different attack hyperparameters (attack iterations, learning rate schedulers, poisoning optimizers), we closely follow the settings from the original papers for each of the attacks in our experiments. We evaluate the attacks with low poison ratios: 5% and 10% of the target-class dataset, equivalent to approximately 0.5% and 1% of the whole dataset. By default, the trigger used in our experiments is sunglasses.
Table 9: Hyperparameters for Gradient Matching attack
@@
In the original Gradient Matching attack, Souri et al. [21] proposed retraining the attacker model and selecting poison samples with maximum gradients to increase ASR. For a fair comparison between poisoning algorithms, however, we do not include such additional factors in our settings. In our experiment, the attacker selects the poison samples randomly.
Table 10: Hyperparameters for Feature Matching attack
Hyperparameters Value
Optimizer Projected Gradient Descend [47]
Learning rate (lr) $0.01$
Attack iterations $5000$
Learning rate scheduler Decay the learning rate by 0.95 every 2000 iterations
Table 11: Hyperparameters for Blended Label Consistent
Hyperparameters Value
Optimizer Projected Gradient Descend
Learning rate (lr) $(\epsilon/4)/225$
Attack iterations $100$
Trigger opacity $32/225$
Learning rate scheduler No scheduler
D.4 Evaluation metrics.
@@
In some cases, CLPBA could be detected if (1) the trigger acts as universal trigger that causes targeted misclassification for any classes, and (2) the model incorrectly recognizes false triggers by unintended objects. Thus, we adopt two new metrics to evaluate the stealthiness of CLPBA:
  • Suspicion Rate (SR): This measures the likelihood that the backdoor attack is unintentionally activated by unintended classes using the physical trigger. It is quantified by the percentage of samples from non-source and non-target classes that are misclassified as the target class when the main trigger is used.
  • False Positive Rate (FPR): This measures the sensitivity of the poisoned model to false triggers. It is quantified by the percentage of samples from the source class that are misclassified as the target class when the false triggers are used.
CLPBA is considered successful if it has a high ASR ($\geq 80\%$) with a negligible decrease ($\leq 5\%$) in ACC. The attack is stealthy if it has a low FPR and a low SR ($\leq 10\%$).
Appendix E Additional experiments
E.1 Experiments on other models
Table 12: Performance of CLPBA ($\alpha=0.1$, $\epsilon=16/226$) on VGG16
Poison pair Naive BLC FM GM
1-7 63.8 (4.4) 2.3 (0.5) 82.3 (12.5) 100.0 (0.0)
3-6 98.6 (1.4) 95.7 (1.4) 100.0 (0.0) 100.0 (0.0)
5-1 42.0 (14.2) 0.3 (0.5) 25.9 (15.8) 86.7 (1.9)
6-8 0.0 (0.0) 0.0 (0.0) 1.4 (1.7) 0.0 (0.0)
7-3 4.0 (3.0) 0.0 (0.0) 89.1 (1.6) 85.2 (7.3)
9-5 0.0 (0.0) 0.0 (0.0) 6.3 (4.0) 6.3 (4.0)
@@ E.1 Experiments on other models
Table 13: Performance of CLPBA ($\alpha=0.1$, $\epsilon=16/226$) on MobileNetV2
Poison pair Naive BLC FM GM
1-7 81.8 (7.1) 3.1 (3.5) 16.0 (11.6) 50.1 (18.3)
3-6 92.3 (6.0) 39.1 (9.1) 51.7 (8.5) 84.5 (5.0)
5-1 87.0 (4.2) 19.4 (7.0) 36.4 (15.2) 43.5 (8.5)
6-8 56.5 (1.3) 24.8 (9.3) 43.3 (1.7) 52.1 (1.7)
7-3 8.3 (8.8) 0.0 (0.0) 0.3 (0.5) 0.0 (0.0)
9-5 4.0 (1.7) 0.0 (0.0) 0.1 (0.2) 15.7 (3.4)
ACC (%) of clean model 93.2 (0.2)
Table 14: Performance of CLPBA ($\alpha=0.1$, $\epsilon=16/226$) on Inception-ResNet
Poison pair Naive BLC FM GM
1-7 5.4 (7.9) 0.3 (0.5) 0.3 (0.5) 96.0 (3.2)
3-6 61.4 (4.7) 48.8 (15.6) 80.2 (1.7) 100.0 (0.0)
5-1 4.3 (1.9) 0.9 (0.0) 2.8 (1.9) 98.5 (1.9)
6-8 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 96.4 (0.5)
7-3 0.3 (0.5) 0.0 (0.0) 0.0 (0.0) 97.8 (2.1)
9-5 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 88.9 (6.9)
@@
The results on MobileNetV2 in Table 13 are different from other models as the Naive attack has a very high ASR compared to other methods. The prediction ACC of MobileNetV2 is also lower with only 93.2%. We believe that MobileNetV2 is less robust than other models, which results in higher ASR with Naive attack. Furthermore, the attack performance on poison pairs can be model dependent as we can see in Table 12 that the pair $6-8$ and $9-5$ fails for VGG16, while poison pair $7-3$ fails for MobileNetV2 (Table 13).
E.2 Scaling the dataset size
@@
We note that since the poison ratio $\alpha$ is defined over the target class dataset, the number of poisoned images does not change when we increase the dataset size. Interestingly, as shown in Table 16 and Table 15, scaling dataset size only has minimal impacts on the attack performance.
E.3 Convergence analysis
We compare the convergence rate of standard $\ell_{\infty}$ regularization with our proposed pixel regularization. Figure 12(b) shows the loss of the GM poisoning objective (Equation 4) and the $\ell_{2}$ norm of the perturbations. Our pixel regularization leads to a better convergence of the poisoning objective while simultaneously reducing the $\ell_{2}$ norm of the perturbations to a smaller value, indicating an improved perceptual quality. Consequently, the GM attack with pixel regularization achieves an ASR of $97.8(1.1)$, compared to $51.9(18.7)$ for the original GM attack under the same conditions.
(a) Standard $\ell_{\infty}$ regularization
(b) Pixel Regularization
Figure 12: Convergence rate of standard $\ell_{\infty}$ regularization and Pixel Regularization. Poisoned samples are crafted with GM attack ($\alpha$ = 0.1, $\epsilon$ = 8/255) on poison pair $7-3$ and ResNet50 model
Appendix F Additional visualization
F.1 Poisoned samples of different poisoning algorithms
@@
F.2 Statistical evaluation
Figure 14: Visualization of poisoned samples of different poisoning algorithms. Perturbations are scaled 10×10$\times$10 × for visibility.
For statistical evaluation, we follow previous work [31, 32] in measuring the Peak Signal-to-Noise Ratio (PSNR) between the poisoned image and the original image. Quantitatively, FM is the most stealthy as the perturbations have the lowest $\ell_{2}$ norm and the poisoned images have the highest PSNR value. Qualitatively, we observe that the noise of FM is more "natural" to the human eye because it is more fine-grained and scattered, resembling natural noise from data augmentation. In comparison, the noise from GM and BLC is smoother with some specific patterns (e.g., in GM, the perturbations are more concentrated on face edges and the area underlying the trigger, while in BLC, the noise has dark areas around the images). We argue that these more uniform, smooth noise patterns from GM and BLC are visually conspicuous to human inspection, compromising the stealthiness of the attacks.
F.3 Additional imperceptibility comparison of regularizations
@@
Appendix G Defenses
To defend against backdoor attacks in DNNs, a multitude of strategies have been developed, each addressing different vulnerabilities. Preprocessing-based defenses [50, 51] aim to weaken embedded triggers by altering the training data, while Model reconstruction-based defenses [52, 53] rectify the poisoned model by eliminating hidden backdoors through retraining, neuron pruning, or knowledge distillation. Trigger synthesis-based defenses [54, 55] counteract backdoors by identifying and neutralizing the specific triggers used in attacks. Model diagnosis-based defenses [56, 57] scrutinize a model’s behavior to detect anomalies indicative of a backdoor. Filtering-based defenses [34, 36, 58, 38, 35] identify and filter suspicious data that are likely to be poisoned by adversaries.
In the paper, we evaluate CLPBA against four representative filtering-based defenses:

### v2

@@ 7 Accidental Backdoor Activations
For CLPBA to be stealthy, we argue that it needs to be source-specific and trigger-specific. That is, (1) only the source class can use the trigger to manipulate the victim model’s prediction at inference time; (2) the trigger should be discriminative so that the victim model does not incorrectly recognize similar-looking objects as triggers. We introduce two new metrics, False Positive Rate, and Suspicion Rate, to test accidental backdoor activations:
@@
Poison pair Trigger Loss type ASR ACC FPR SR
9-5 Fake beard Original 98.7 99.9 50.8 12.5
With repel 98.7 99.9 49.7 3.9
7-3 Sunglasses Original 97.5 99.3 0.7 13.9
With repel 97.5 99.4 0.9 2.5
Table 5: FPR and SR for GM attack. The set-up of our dataset to evaluate SR and FPR is given in Appendix B.
@@
$L_{adv}=N\cdot\overbrace{\sum_{x\in\widetilde{D}_{s}}\mathcal{L}\bigl{(}F_{\theta}(x),y_{t}\bigr{)}}^{\text{adversarial loss}}$
[…]
@@ 8 Conclusion
Through extensive experiments, our research underscores the vulnerability of DNNs to clean-label physical backdoor attacks. We then show that CLPBA works by embedding adversarial features and trigger features in the poisoned instances. Additionally, we propose a novel pixel regularization to improve the effectiveness and stealthiness of poisoned samples. A significant limitation identified is accidental backdoor activations of the backdoored model due to false triggers or non-source classes, which could compromise attack stealthiness. To address this, we propose incorporating repel terms into the adversarial objective to enhance trigger specificity to the source class. Our findings highlight the critical need for robust defenses against CLPBA. By releasing the dataset and code, we hope to facilitate future research on improving the resilience of DNNs against such threats.
Supplementary Material
Appendix A Natural Physical Backdoors
@@
(a) Clean samples (w/o trigger)
@@
Appendix B Dataset
B.1 Ethics & Data Protection
IRB approval. Before conducting our study, we submitted a ”Human Subjects Research – Expedited Review Form” to our country’s Institutional Review Board (IRB). Our study received approval from the chairman of the institutional ethical review board under decisions No 24/2016/QD-VINMEC, No 23/2016/QD-VINMEC, and No 77/2021/QD-VINMEC. We prepared a consent form beforehand to ensure transparency in the procedure of dataset collection. All 10 volunteers in our dataset provided explicit written consent for us to collect the dataset and use the images for research purposes, including permission to use the captured images in the research paper.
@@ B.4 Dataset Structure
False triggers. To evaluate the sensitivity of poisoned models to false triggers, as measured by false positive rate (FPR) (see Appendix C.4), we select three additional false triggers for each of the main triggers. Because it is hard to find a representative set of false triggers for each of the main triggers, we select the false triggers by changing the sizes, colors, and shapes of the objects (e.g, we change the shape of earrings and hats, change the colors of the headband, change the size of sticker). We also select objects that have similar locations on the face as false triggers (e.g., headbands are selected as false triggers for the hat trigger and the sunglasses trigger). In Figure 17, each row corresponds to a main trigger on the left, followed by three false triggers on the right.
@@
B.5 Animal Classification Dataset
Besides facial recognition, we also conduct an extended evaluation of CLPBA on animal classification. We collected 2000 cat images with 500 trigger-free images and 500 images for each of the triggers, which are tennis balls, mobile phones, and books. The trigger-free cat images are resized to $224\times 224$ and then concatenated to an existing animal classification dataset on Kaggle. Visualization of trigger images for this dataset is given in Figure 16  [^1] https://www.kaggle.com/datasets/asaniczka/mammals-image-classification-dataset-45-animals.
Appendix C Experiment Settings
C.1 Computational Resources.
All the experiments are conducted on the two servers. The first one with 7 RTX 3090 Ti 24 GB GPUS, and the second one contains 6 RTX A5000 24 GB GPUS. The code is implemented in Pytorch. We develop the framework based on the previous works [49, 15]. We use Pytorch as the framework for our model training.
C.2 Training Settings.
We evaluate the attacks on four Convolutional Neural Networks for facial recognition: ResNet50 [18], VGG16 [48], MobileNetV4 [45], Inception-ResNet [50]. We choose fine-tuning as the training method for the victim model since it is a widely adopted training method for its speed of convergence and accuracy. Compared to transfer-learning, Fine-tuning and training from scratch are shown to be hard scenarios for clean-label attacks [44, 49, 15]. Our models are pre-trained on VGGFace2 [4], a large-scale facial dataset containing 3.31 million images divided into 9131 classes. Since BLC requires an adversarially trained model, we perform fast adversarial training [57] for the pre-trained model on our dataset before crafting poisons. We do not apply any data augmentations to the training pipeline. We only standardize the training data using z-score normalization with the mean and standard variation calculated from the data. We observe that the training models converge very quickly and achieve the validation accuracy of over $99\%$ and $93\%$ by epoch 10 for Facial Recognition and Animal Classification tasks.
Table 8: Training Hyperparameters
@@ Appendix C Experiment Settings
Table 9: Hyperparameters for Poisoning Algorithms
@@
Trigger opacity (BLC) $32/225$
Inner SGD steps (MP) $2$
C.3 Attack Settings.
Our experiments follow the gray-box scenario in which the attacker model has the same architecture and is trained on the same dataset as the victim model but has different weight initialization, and the attacker cannot access and modify the weights and training procedure of the victim model. The dataset of trigger samples is split 50-50 into train and test sets. The train set is used for crafting poisons while the test set is used to validate the attacks. Table 9 gives the hyperparameters for poisoning algorithms.
In the original Gradient Matching attack, Souri et al. [49] proposed retraining the attacker model and selecting poison samples with maximum gradients to increase ASR. For a fair comparison between poisoning algorithms, however, we do not include such factors in our settings. In our experiment, the attacker selects the poison samples randomly.
C.4 Evaluation Metrics.
@@
Appendix D Additional Experiments
D.1 Blended Label-Consistent Attack
For completeness, in our evaluation, we also evaluate Blended Label-Consistent attack [53, 8] which works by perturbing the poisoned samples with adversarial noise to make the salient features of the samples harder to learn before injecting the digital trigger, forcing the model’s classification to depend on the trigger.
Following [53], we first train an adversarial-trained model, before using the poisoning method in [53] to craft adversarial examples. Then we insert the digital trigger $p$ (e.g., digital sunglasses) via a perspective transformation $T$ that maps the trigger to the appropriate facial area using facial landmarks. The trigger is inserted with an opacity $\tau$.
$x_{t}$ $\leftarrow(1-\tau)\cdot x_{t}+\tau\cdot T(p)$ (15)
The procedure of BLC is illustrated in Figure 11. Compared to random placement and fixed center placement of the trigger, our adaptive trigger placement method with facial landmarks has a higher attack success rate as it preserves the spatial relationship between the trigger and the face.
D.2 Experiments on Different Models
Table 10: CLPBA ($\alpha=0.1$, $\epsilon=16/226$) on VGG16
Poison pair Naive BLC MP FM GM
1-7 63.8 (4.4) 2.3 (0.5) 0.3 (0.5) 82.3 (12.5) 100.0 (0.0)
3-6 98.6 (1.4) 95.7 (1.4) 39.6 (9.9) 100.0 (0.0) 100.0 (0.0)
5-1 42.0 (14.2) 0.3 (0.5) 0.9 (0.0) 25.9 (15.8) 86.7 (1.9)
6-8 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.4 (1.7) 0.0 (0.0)
7-3 4.0 (3.0) 0.0 (0.0) 0.0 (0.0) 89.1 (1.6) 85.2 (7.3)
9-5 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 6.3 (4.0) 6.3 (4.0)
@@ D.2 Experiments on Different Models
Table 11: CLPBA ($\alpha=0.1$, $\epsilon=16/226$) on Inception-ResNet
Poison pair Naive BLC MP FM GM
1-7 5.4 (7.9) 0.3 (0.5) 13.7 (9.9) 0.3 (0.5) 96.0 (3.2)
3-6 61.4 (4.7) 48.8 (15.6) 96.1 (2.2) 80.2 (1.7) 100.0 (0.0)
5-1 4.3 (1.9) 0.9 (0.0) 0.0 (0.0) 2.8 (1.9) 98.5 (1.9)
6-8 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 96.4 (0.5)
7-3 0.3 (0.5) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 97.8 (2.1)
9-5 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 88.9 (6.9)
@@
We further evaluate the performance of CLPBA on 6 random poison pairs, on VGG16 and Inception-ResNet model architecture for the Facial Recognition task in Table 10 and Table 11.
D.3 Scaling the Dataset Size
@@
Since the poison ratio $\alpha$ is defined over the target class dataset, the number of poisoned images does not change when we increase the dataset size. Interestingly, as shown in Table 13 and Table 12, scaling dataset size only minimally impacts the attack performance.
D.4 Evaluating Digital Trigger for CLPBA
In CLPBA, we assume that the attackers hold trigger samples from the source class to perform poisoning. However, this may not be practical in real-world scenarios when the attacker can access clean images from the source class but not the trigger images. To mitigate this, the attacker can inject a digital image of the physical trigger into the source-class images as an alternative. In this section, we aim to evaluate the performance of CLPBA when the poisoned samples are crafted with photoshopped trigger images. As in Table 14, using photoshopped images has varying effectiveness for different poison pairs.
Table 14: Evaluating the use of digitally-modified trigger images for GM attack. The metric reported is ASR (%)
Poison pairs 1-7 2-0 6-8 9-5
GM-digital 98.0 (1.0) 0.0 (0.0) 68.6 (8.6) 66.3 (21.5)
GM-physical 100.0 (0.0) 99.4 (1.0) 98.3 (0.8) 97.8 (0.2)
D.5 Evaluating transferability of CLPBA
We evaluate the black-box settings of CLPBA for two poison pairs $7-1$ and $3-6$ in Table 15 and Table 16 Generally, we find that CLPBA under black-box settings have a worse performance, which means that the semantic features extracted by different model architectures may differ greatly. The same attack on different poison pairs may have vastly different results. For example, the poison pair $3-6$ with sunglasses trigger has a much better transferability than the poison pair $7-1$. This shows that the effectiveness of the trigger depends on the alignment between the trigger and natural features of the target class in the physical space.
Table 15: Evaluate transferability of CLPBA on poison pair $1-7$
Attacker Victim Inception-ResNetV1 ResNet50 VGG16
GM FM GM FM GM FM
Inception-ResNetV1 96.9 0.3 46.4 16.2 20.8 5.4
ResNet50 1.7 0.3 99.7 92.0 35.3 6.3
VGG16 3.7 0.3 91.5 24.2 100.0 83.2
Table 16: Evaluate transferability of CLPBA on poison pair $3-6$
Attacker Victim Inception-ResNetV1 ResNet50 VGG16
GM FM GM FM GM FM
Inception-ResNetV1 100.0 79.2 30.4 18.8 99.5 97.6
ResNet50 72.9 55.6 100.0 92.3 98.6 98.1
VGG16 72.9 49.3 71.5 22.2 100.0 100.0
Appendix E Additional Visualization
E.1 Poisoned samples of different poisoning algorithms
@@
E.2 Additional imperceptibility comparison of regularizations
@@
Appendix F Defenses
To defend against backdoor attacks in DNNs, a multitude of strategies have been developed, each addressing different vulnerabilities. Preprocessing-based defenses [33, 10] aim to weaken embedded triggers by altering the training data, while Model reconstruction-based defenses [32, 62] rectify the poisoned model by eliminating hidden backdoors through retraining, neuron pruning, or knowledge distillation. Trigger synthesis-based defenses [54, 7] counteract backdoors by identifying and neutralizing the specific triggers used in attacks. Model diagnosis-based defenses [22, 58] scrutinize a model’s behavior to detect anomalies indicative of a backdoor. Filtering-based defenses [52, 3, 14, 40, 17] identify and filter suspicious data that are likely to be poisoned by adversaries. Firewall-based defenses aim to detect malicious inputs at inference time and discard them so that the trained backdoored model will not behave maliciously.
In the paper, we evaluate CLPBA against four representative filtering-based defenses and four representative firewall defenses: